A spherical metal shell A of radius RA and a solid metal sphere B of radius RB (A) are kept far apart and each is given charge ‘+Q’, Now they are connected by a thin metal wire Then
 
(a) EAinside = 0
(b) QA> QB
(c) σ A/ σB = RB/RA
(d) EAonsurfaceBonsurface

Grade:11

1 Answers

Kevin Nash
9 years ago
Hello Student,
Please find the answer to your question
. (a,b,c,d)
Electric field inside a spherical metallic shell with charge on the surfaceis always zero. Therefore option [a] is
correct.
When the sells are connected with a thin metal wire then electric potentials will be equal, say V
∴ 1/ 4πϵ0 QA/RA = 1 / 4 πϵ0 QB­/RB = V
As RA> RB therefore QA> A­B option [b] is also correct.
As σA/σB QA/4πR2A/QB/4πR2B = R2B / R2A x 4πϵ0 RAV / 4πϵ0 RAV
∴σAB = RB / RA
Option (c) is also correct
Also EA = σA/ σB = RB/RA< 1 ∴ EA< EB
Option (d) is also correct
Thanks
Kevin Nash
askIITians Faculty